# Why does the Milky Way have a metallicity floor?

### v1

@@ Abstract
The abundance patterns of extremely metal-pool stars comprise nearly all of the observational data that we have on the process of star formation below a metallicity of about 10-3 Z⊙. One of the most notable features of these objects, collectively, is their enhancement in light elements, namely C and O, with respect to Fe, the classical proxy for “metallicity.” From this, we have the notion of Carbon-Enhanced Metal-Poor (CEMP) stars (Beers & Christlieb, 2005), which comprise a majority of stars with [Fe/H] -3. If we consider the combined metal content of these stars, however, another interpretation of the CEMP phenomenon is that of a metallicity floor (e.g., Starkenburg et al., 2018). Indeed, as noted there and elsewhere, [C/Fe] seems to increase with decreasing [Fe/H] for the most metal-poor objects. This point was made succinctly by Frebel et al. (2007), who introduce a combined C/O abundance measure known as the transition discriminant (), given by
@@
When viewed from this perspective (e.g., Figure 1 of Frebel et al., 2007), the Galaxy appears to have a metallicity floor of , equivalent to an absolute metallicity of about 10-3.6 Z⊙.
@@
This number has very interesting astrophysical significance. Well above this metallicity, observations overwhelmingly support the existence of a roughly universal stellar initial mass function (IMF), robust to physical conditions and producing predominantly low mass stars (see, e.g., Krumholz, 2014b)111In practice, direct observational evidence of the IMF extends down to around 10-2 Z⊙ (see the discussion in Sharda & Krumholz, 2022).. Well below it (specifically, zero metallicity), we expect the star formation process to have fundamentally different results, i.e., a distinctly top-heavy IMF that may also have environmental dependence (see, e.g., Klessen & Glover, 2023). To zeroth order we know that H2 (the main coolant in cold, metal-free gas) plants the initial seed for a top-heavy IMF through its inefficient cooling, which limits Jeans fragmentation to mass scales of 1000 M⊙222We note that most simulations now show the dense accretion disk surrounding the massive pre-stellar core will readily fragment into small clumps (again, see Klessen & Glover, 2023), but this does not alter the main conclusion.. If we then consider the point during the collapse of metal-free gas where H2 cooling becomes inefficient and fragmentation ends, it is possible to calculate the amount of metal required for continued cooling and fragmentation, arriving at a metallicity of 10-3.5 Z⊙. Specifically, this is done by equating the cooling time with the free-fall time at the relevant density and temperature ( cm-3 and K). What’s more, as was first pointed out by Bromm & Loeb (2003), the most important metal coolants in this regime come from atomic C and O. Thus, it is tempting to conclude that the Galaxy’s metallicity floor traces the “gas-phase critical metallicity” for the transition from Population (Pop) III to Pop II star formation.
@@
Problems exist with this interpretation, however. Another significantly lower critical metallicity for fragmentation exists due to cooling from dust (Omukai et al., 2005), with a value of 10-5.5 Z⊙ (Chiaki et al., 2015, assuming a local ISM dust-to-metal ratio). Decoupling dust from gas-phase metallicity, this value has also been expressed as a critical “dusticity,” (Schneider et al., 2012)333This is about 5 of the local dust-to-gas ratio of 0.009387 (Pollack et al., 1994)., still far less total metal than the previously stated gas-phase critical metallicity. As well, both observations and simulations indicate that the Universe has the ability to create low mass stars above the dust-phase critical metallicity but below the gas-phase one. Famously, the star SDSS J102915+172927 (Caffau et al., 2011) has [Fe/H] -5 and [C/H] -4.4. Complementing that, the fully cosmological simulations of Smith et al. (2015) showed that such a low metallicity (in that work, Z⊙) star-forming environment could arise in a minihalo on the cusp of Pop III star formation that is enriched by a supernova blast-wave originating from another nearby minihalo, i.e., prompt external enrichment. In summary, the Galactic metallicity floor does not appear to be a signal of IMF transition.
@@ 3.2.1 Model Summary
As we wish to model the gas in each of our simulated minihalos individually, we must first construct the time-dependent data characterizing the influences experienced as the halos grow and other stars form nearby. We create radial profiles of each halo from every simulation snapshot in which it exists and remap them from radius to enclosed gas mass (e.g., Figure 1). We identify the mass coordinate with the highest ratio of enclosed mass to Bonnor-Ebert mass in the snapshot immediately prior to Pop III star particle formation. We then initialize the relevant gas quantities (i.e., density, internal energy, chemical species fractions) using the values in that mass coordinate at the time when the halo’s most massive progenitor has a virial mass of roughly 104 M⊙. We also make note of the radius corresponding to the target mass coordinate in the profile. This serves as the initial radius of the gas parcel, which we update as the density changes, assuming spherical symmetry. As the model runs we update the dark matter density (used in the modified free-fall component) and photo-chemical/heating rates (stimuli to the chemistry solver) using the radial profiles at the relevant cosmic time and mass coordinate. From one time step to the next, we update the internal energy and chemical species using the Grackle444https://grackle.readthedocs.io/ chemistry and cooling library (Smith et al., 2017) with the same settings as were used in the simulation.
@@ 3.2.3 The Importance of Gravity, Pressure, and Radiation
Regardless, it is clear that thermal pressure is important for regulating the rate of collapse. In their one-zone models, Omukai et al. (2005) account for this by modifying the free-fall collapse model introduced in Omukai (2000) to slow the density evolution by a factor relating to the ratio of pressure gradient forces to gravity (i.e., Equations 3, 5–7), and setting a maximum value of . For the halos in this work this generally results in collapse proceeding too quickly, as illustrated by the dashed line in Figure 3. Increasing the upper limit on to 0.99 further slows collapse, but in the case of halo 10/11 (Figure 3, long dashed line) by too much, whereas in other cases it is still not enough. It is also clear from Figure 3 that this approximation fails to reproduce the low density, pressure-dominated evolution entirely555It should be noted that, while upper limits on greater than 0.95 resulted in collapse times both too short and too long, a value of 0.95 was fairly consistently too short and likely the best value to use..
@@ Acknowledgements
BDS wishes to thank Heather Jacobson, John Regan, and the entire FOGGIE collaboration (in particular, Molly Peeples) for useful conversations and continued support. BDS and SK are supported by STFC Consolidated Grant RA5496. BWO acknowledges support from NSF grants #1908109 and #2106575 and NASA ATP grants NNX15AP39G and 80NSSC18K1105. JHW acknowledges support from NSF grant #2108020 and NASA grant 80NSSC20K0520. BWO and JHW acknowledge support from NASA TCAN grant 80NSSC21K1053. This research is part of the Blue Waters sustained-Petascale Computing Project, which is supported by the NSF (award number ACI-1238993) and the state of Illinois. Blue Waters is a joint effort of the University of Illinois at Urbana-Champaign and the National Center for Supercomputing Applications. The simulations described by this paper were run using an NSF PRAC allocation (award number OCI-0832662). For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
@@
All codes to produce the analysis and figures for this work are available in the yt_p2p666https://github.com/brittonsmith/yt_p2p yt extension and the “p2p” branch of the lead author’s fork of Grackle777https://github.com/brittonsmith/grackle/tree/p2p. Data products used in this work include simulation snapshots, halo catalogs, merger trees, and halo radial profiles. These will be made available upon reasonable request to the lead author.

### v2

@@ Abstract
The abundance patterns of extremely metal-pool stars comprise nearly all of the observational data that we have on the process of star formation below a metallicity of about 10-3 Z⊙. One of the most notable features of these objects, collectively, is their enhancement in light elements, namely C and O, with respect to Fe, the classical observational proxy for “metallicity111While iron abundance, [Fe/H], has often been used historically to denote metallicity, we make a distinction in this work between individual elemental abundances and metallicity. We use metallicity, , to mean the fraction of total gas mass in metals, whereas elemental abundances are typically expressed as the ratio of two elements, X and Y, by [X/Y] , where is the number abundance of element X.” From this, we have the notion of Carbon-Enhanced Metal-Poor (CEMP) stars (Beers & Christlieb, 2005), which comprise a majority of stars with [Fe/H] -3. If we consider the combined metal content of these stars, however, another interpretation of the CEMP phenomenon is that of a metallicity floor (e.g., Starkenburg et al., 2018). Indeed, as noted there and elsewhere, [C/Fe] seems to increase with decreasing [Fe/H] for the most metal-poor objects. This point was made succinctly by Frebel et al. (2007), who introduce a combined C/O abundance measure known as the transition discriminant (), given by
@@
This particular abundance measure is designed to highlight the most important coolants in low metallicity gas (see below and Section 4). Frebel et al. (2007) find a minimum value for stars in the Milky Way of (Figure 1 of that work), equivalent to an absolute metallicity of about 10-3.6 Z⊙. Thus, when viewed from this perspective, the Galaxy appears to have a metallicity floor.
@@
This number has very interesting astrophysical significance. Well above this metallicity, observations overwhelmingly support the existence of a roughly universal stellar initial mass function (IMF), robust to physical conditions and producing predominantly low mass stars (see, e.g., Krumholz, 2014b)222In practice, direct observational evidence of the IMF extends down to around 10-2 Z⊙ (see the discussion in Sharda & Krumholz, 2022).. Well below it (specifically, zero metallicity), we expect the star formation process to have fundamentally different results, i.e., a distinctly top-heavy IMF that may also have environmental dependence (see, e.g., Klessen & Glover, 2023). To zeroth order we know that H2 (the main coolant in cold, metal-free gas) plants the initial seed for a top-heavy IMF through its inefficient cooling, which limits Jeans fragmentation to mass scales of 1000 M⊙333We note that most simulations now show the dense accretion disk surrounding the massive pre-stellar core will readily fragment into small clumps (again, see Klessen & Glover, 2023), but this does not alter the main conclusion.. If we then consider the point during the collapse of metal-free gas where H2 cooling becomes inefficient and fragmentation ends, it is possible to calculate the amount of metal required for continued cooling and fragmentation, arriving at a metallicity of 10-3.5 Z⊙. Specifically, this is done by equating the cooling time with the free-fall time at the relevant density and temperature ( cm-3 and K). What’s more, as was first pointed out by Bromm & Loeb (2003), the most important metal coolants in this regime come from atomic C and O. Thus, it is tempting to conclude that the Galaxy’s metallicity floor traces the “gas-phase critical metallicity” for the transition from Population (Pop) III to Pop II star formation.
@@
Problems exist with this interpretation, however. Another significantly lower critical metallicity for fragmentation exists due to cooling from dust (Omukai et al., 2005), with a value of 10-5.5 Z⊙ (Chiaki et al., 2015, assuming a local ISM dust-to-metal ratio). Decoupling dust from gas-phase metallicity, this value has also been expressed as a critical “dusticity,” (Schneider et al., 2012)444This is about 5 of the local dust-to-gas ratio of 0.009387 (Pollack et al., 1994)., still far less total metal than the previously stated gas-phase critical metallicity. As well, both observations and simulations indicate that the Universe has the ability to create low mass stars above the dust-phase critical metallicity but below the gas-phase one. Famously, the star SDSS J102915+172927 (Caffau et al., 2011) has [Fe/H] -5 and [C/H] -4.4. Complementing that, the fully cosmological simulations of Smith et al. (2015) showed that such a low metallicity (in that work, Z⊙) star-forming environment could arise in a minihalo on the cusp of Pop III star formation that is enriched by a supernova blast-wave originating from another nearby minihalo, i.e., prompt external enrichment. In summary, the Galactic metallicity floor does not appear to be a signal of IMF transition.
@@ 3.2.1 Model Summary
As we wish to model the gas in each of our simulated minihalos individually, we must first construct the time-dependent data characterizing the influences experienced as the halos grow and other stars form nearby. We create radial profiles of each halo from every simulation snapshot in which it exists and remap them from radius to enclosed gas mass (e.g., Figure 1). We identify the mass coordinate with the highest ratio of enclosed mass to Bonnor-Ebert mass in the snapshot immediately prior to Pop III star particle formation. We then initialize the relevant gas quantities (i.e., density, internal energy, chemical species fractions) using the values in that mass coordinate at the time when the halo’s most massive progenitor has a virial mass of roughly 104 M⊙. We also make note of the radius corresponding to the target mass coordinate in the profile. This serves as the initial radius of the gas parcel, which we update as the density changes, assuming spherical symmetry. As the model runs we update the dark matter density (used in the modified free-fall component) and photo-chemical/heating rates (stimuli to the chemistry solver) using the radial profiles at the relevant cosmic time and mass coordinate. From one time step to the next, we update the internal energy and chemical species using the Grackle555https://grackle.readthedocs.io/ chemistry and cooling library (Smith et al., 2017) with the same settings as were used in the simulation.
@@ 3.2.3 The Importance of Gravity, Pressure, and Radiation
Regardless, it is clear that thermal pressure is important for regulating the rate of collapse. In their one-zone models, Omukai et al. (2005) account for this by modifying the free-fall collapse model introduced in Omukai (2000) to slow the density evolution by a factor relating to the ratio of pressure gradient forces to gravity (i.e., Equations 3, 5–7), and setting a maximum value of . For the halos in this work this generally results in collapse proceeding too quickly, as illustrated by the dashed line in Figure 3. Increasing the upper limit on to 0.99 further slows collapse, but in the case of halo 10/11 (Figure 3, long dashed line) by too much, whereas in other cases it is still not enough. It is also clear from Figure 3 that this approximation fails to reproduce the low density, pressure-dominated evolution entirely666It should be noted that, while upper limits on greater than 0.95 resulted in collapse times both too short and too long, a value of 0.95 was fairly consistently too short and likely the best value to use..
@@ Acknowledgements
We thank the referee, Gerry Gilmore, for helpful comments on the manuscript. BDS wishes to thank Heather Jacobson, John Regan, and the entire FOGGIE collaboration (in particular, Molly Peeples) for useful conversations and continued support. BDS and SK are supported by STFC Consolidated Grant RA5496. BWO acknowledges support from NSF grants #1908109 and #2106575 and NASA ATP grants NNX15AP39G and 80NSSC18K1105. JHW acknowledges support from NSF grant #2108020 and NASA grant 80NSSC20K0520. BWO and JHW acknowledge support from NASA TCAN grant 80NSSC21K1053. This research is part of the Blue Waters sustained-Petascale Computing Project, which is supported by the NSF (award number ACI-1238993) and the state of Illinois. Blue Waters is a joint effort of the University of Illinois at Urbana-Champaign and the National Center for Supercomputing Applications. The simulations described by this paper were run using an NSF PRAC allocation (award number OCI-0832662). For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
@@
All codes to produce the analysis and figures for this work are available in the yt_p2p777https://github.com/brittonsmith/yt_p2p yt extension and the “p2p” branch of the lead author’s fork of Grackle888https://github.com/brittonsmith/grackle/tree/p2p. Data products used in this work include simulation snapshots, halo catalogs, merger trees, and halo radial profiles. These will be made available upon reasonable request to the lead author.